OpenStudy (anonymous):

Any assistance would be greatly appreciated. Write a system of two equations in two variables to solve the problem. It cost a family of 6 a total of $182 for general admission tickets to the San Diego Zoo. How many adult tickets and how many child tickets were purchased? ________ adult tickets ________ child tickets

OpenStudy (anonymous):

what is the price for each type of ticket?

OpenStudy (texaschic101):

a = 6 - c or c = 6 - a...but I do not know what to do after this

OpenStudy (133):

Not Enough Information Given there is no saying how much the adult tickets are or the children tickets so there is your answer there except for NEIG!!!!!! Hope this helps!!! lol

OpenStudy (texaschic101):

I agree..not enough information....kinda need to know how much adult tickets and how much child tickets cost

OpenStudy (133):

No it says how many were purchased!!!!

OpenStudy (anonymous):

Ah sorry, Adults $37, child $27

OpenStudy (anonymous):



OpenStudy (anonymous):

dalia_lam, Thank you! When comparing, I see where I went wrong too - thank you so much for laying it out since that is where I found the real answer to get it right.

OpenStudy (arthur_ser):

Let x the number of adults and y the number of children in the family respectively. Since the number of the members of the family equlals 6: x + y = 6 and because adults' tickets cost 37$ and chidren's cost 27$ each: 37x + 27y = 182 so you have to solve the following system of equations: x + y =6 37x + 27y =182

OpenStudy (133):

Sorry wasn't here earlier!!!!!!!! LOL :P

OpenStudy (anonymous):

133 - Thanks~!!